To determine the concentration of a solution of nitric acid, a 150.0-mL sample is placed in a flask and titrated with a 0.1175 M solution of barium hydroxide. A volume of 31.46 mL is required to reach the phenolphthalein endpoint. Calculate the concentration of nitric acid in the original sample.

_______ M
1
Expert's answer
2019-03-11T08:19:50-0400

The balanced equation of this reaction:



The total amount of moles of barium hydroxide is:


"n(Ba(OH)_2)= C(Ba(OH)_2)*V(Ba(OH)_2)=0.1175M*0.03146L=0.0037 moles"

According to stoichoiometric quotients, there are 2 moles of nitric acid per each mole of barium hydroxide, so we get:


"n(HNO_3)=2n(Ba(OH)_2)=0.0037*2 moles= 0.0074 moles"

According to molar concentration formula:

"C_M(HNO_3)=\frac n V=0.0074 moles/0.15L = 0.0493 M"

Answer: the concentration of the nitric acid in the sample is 0.0493M.